A) A buffer solution contains 0.373 M NaH2PO4 and 0.348 M Na2HPO4. Determine the pH change when 0.083 mol HBr is added to 1.00 L of the buffer.

pH change = ___

B) A buffer solution contains 0.330 M NH4Br and 0.379 M NH3 (ammonia). Determine the pH change when 0.086 mol HClO4is added to 1.00 L of the buffer.

pH after addition − pH before addition = pH change = ___
Answer #1

A)

[NaH2PO4] = 0.373M

[Na2HPO4] = 0.348M

Volume = 1.00L

number of moles of NaH2PO4= 0.373Mx1.0L= 0.373 moles

number of moles of Na2HPO4= 0.348Mx1.0L= 0.348 moles

Ka of H2PO4-=6.3x10^-8

-log(Ka) = -log(6.3x10^-8)

PKa= 7.2

PH= PKa+log[salt/acid]

PH= 7.2+log( 0.348/0.373)

PH=7.17

number of moles of HBr = 0.083 moles

after addition of HBr

Number of moles of NaH2PO4= 0.373+0.083=0.456 moles

number of moles of Na2HPO4= 0.348 - 0.083=0.265 moles

PH= 7.20+log(0.265/0.456)

PH=6.96

PH change = 6.96 - 7.17 = - 0.21

PH change = -0.21

B)

[NH4Br] = 0.330M

[NH3] = 0.379M

Volume = 1.0L

number of moles of NH4Br= 0.330Mx1.0L= 0.330 moles

number of moles of NH3 = 0.379Mx1.0L= 0.379 moles

PKb of Nh3= 4.75

POH= 4.75+log(0.330/0.379)

POH= 4.69

PH+POH=14

PH= 14-POH

PH= 14-4.69

PH= 9.31

number of moles of HClO4= 0.086 moles

after addition of HClO4

number of moles of NH4Cl = 0.330+0.086=0.416 moles

number of mole sof NH3= 0.379 - 0.086=0.293 moles

POH= 4.75+log(0.416/0.293)

POH=4.90

PH= 14-4.90

PH=9.1

PH change = 9.1 - 9.31= - 0.21

PH change = - 0.21.